OpenStudy (anonymous):

show sin^2xcos^2x=(1-cos4x)/8

OpenStudy (anonymous):

Trig identites

OpenStudy (experimentx):

sin^2xcos^2x = 1/4 (4 sin^2x cos^2x) = (2sin x cos x)^2 / 4 = (sin2x)^2 / 4 = (1 - 2cos4x)/ 2*4 = (1 - 2cos4x)/ 8